An electron is accelerated by a constant electric field of magnitude 300 N/C.
(a) Find the acceleration of the electron.
1Your answer was incorrect, but has changed from what was graded.
Your response differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. m/s2

(b) Use the equations of motion with constant acceleration to find the electron's speed after 8.00 10-9 s, assuming it starts from rest.
2 m/s
Answer #1

Given

\(E=300 \mathrm{~N} / \mathrm{C}\)

\(e=1.60 \times 10^{-10} \mathrm{C}\)

\(m=9.11 \times 10^{-31} \mathrm{~kg}\)

\((a)\) acceleration \(a=\frac{E e}{m}\)

\(=0.527 \times 10^{14} \mathrm{~m} / \mathrm{s}^{2}\)

\(=527 \times 10^{11} \mathrm{~m} / \mathrm{s}^{2}\)

(b)

Given

\(\begin{aligned} u &=0 \\ t=& 8.00 \times 10^{-9} \mathrm{~s} \\ v &=u+a t \\ &=0+\left(527 \times 10^{11} \mathrm{~m} / \mathrm{s}^{2}\right)\left(8.00 \times 10^{-9} \mathrm{~s}\right) \\ &=4216 \times 10^{2} \mathrm{~m} / \mathrm{s} \\ &=421.6 \mathrm{Km} / \mathrm{s} \\ & \approx 422 \mathrm{~km} / \mathrm{s} \end{aligned}\)